# Statistical learnability of smooth boundaries via pairwise binary classification with deep ReLU networks

### v1

@@ 2.2 Assumptions
We next recall the classes of smooth boundaries introduced by Imaizumi and Fukumizu, (2022). For a given $\alpha>0$ and $d\in\mathbb{N}$, denote by $\mathcal{C}^{\alpha,d}$, the unit ball of the $\alpha$-Hölder space on $[0,1]^{d}$, namely
$\mathcal{C}^{\alpha,d}=\left\{g:[0,1]^{d}\to\mathbb{R}\;|\;g\textup{ is }\alpha_{*}\textup{-times differentiable, }\|g\|_{\mathcal{C}^{\alpha,d}}\leq 1\right\},$
@@
    Given $\mathscr{S}=\{\mathcal{J}_{i}\}_{i=1}^{d_{1}}$, there are disjoint subsets $\mathcal{I}_{1},\cdots,\mathcal{I}_{d_{1}}\subset\{1,-1\}^{d_{*}}$, functions $h_{1},\cdots,h_{d_{*}}\in\mathcal{C}^{\alpha,d_{0}-1}$, and indices $j_{1},\cdots,j_{d_{*}}\in\{1,\cdots,d_{0}\}$, such that for every $i=1,\cdots,d_{1}$,
@@
    Given $\mathscr{S}^{\prime}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$, there are disjoint subsets $\mathcal{I}_{1},\cdots,\mathcal{I}_{d_{1}}\subset\{1,-1\}^{d_{*}}$, functions $h_{1},\cdots,h_{d_{*}}\in\mathcal{C}^{\alpha,d_{0}-1}$, and indices $j_{1},\cdots,j_{d_{*}}\in\{1,\cdots,d_{0}\}$ such that for every $i=1,\cdots,d_{1}$,
@@
Since any partition $\mathscr{S}$ in $\mathscr{P}_{\alpha}$ is expressed as a sequence, we need to introduce an equivalence relation that identifies the sequences containing the same subsets: namely, given two sequences $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ and $\{\mathcal{K}_{i}^{\prime}\}_{i=1}^{d_{1}}$ of subsets in $\mathcal{X}$, we define the equivalence relation $\sim$ as $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\sim\{\mathcal{K}_{i}^{\prime}\}_{i=1}^{d_{1}}$ if and only if there is a permutation $\pi$ of the set $\{1,\cdots,d_{1}\}$ such that $\mathcal{K}_{\pi(i)}=\mathcal{K}_{i}^{\prime}$ for any $i\in\{1,\cdots,d_{1}\}$. Denote the equivalence class where the sequence $\mathscr{S}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ belongs to, by $[\mathscr{S}]$.
@@
Given $\alpha>0$ and $P\in\mathcal{P}_{\alpha}$, let $\mathscr{S}\in\mathscr{P}_{\alpha}$ be a partition satisfying the condition (A4) in Definition 1. Then, we define as $\mathscr{S}_{P}=[\mathscr{S}]$.
Note that by the condition (4), the map $P\mapsto\mathscr{S}_{P}$ on $\mathcal{P}_{\alpha}$ is well-defined. Also, given an equivalence class $[\mathscr{S}]\in\mathscr{P}_{\alpha}/\sim$, without loss of generality we can fix an arbitrary partition $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\in[\mathscr{S}]$ in the proof of the main results of this work, by the definition of the equivalence relation.
@@ 2.3 Formal Statement of the Problem
Here, $\|\cdot\|_{1}$ denotes the 1-norm in the Euclidean space. Given $\alpha>0$, let $P\in\mathcal{P}_{\alpha}$. Then, we aim at estimating the sets in the sequence $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\in\mathscr{S}_{P}$. Similarly to Imaizumi and Fukumizu, (2019, 2022), we analyze the convergence rates of estimators in terms of the $L^{2}$-risk. Formally, the estimation error of an estimator $\widehat{g}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$ is measured by the sum of $L^{2}$-risks
@@
Given $P\in\mathcal{P}_{\alpha}$, is there an estimator $\widehat{g}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$ such that the $L^{2}$-risk (5) defined with any sequence $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ in $\mathscr{S}_{P}$ tends to zero as $n\to\infty$? In particular:
@@
    Is it reasonable to use the pairwise dependency to estimate smooth boundaries?
@@ 2.4 Validity of the Estimation Problem
For every $\mathscr{S}\in\mathscr{P}_{\alpha,+}$, there is a Borel probability measure $P\in\mathcal{P}_{\alpha}$ such that we have $[\mathscr{S}]=\mathscr{S}_{P}$.
@@ 3.1 Preliminary
Let $P\in\mathcal{P}$. The contrastive function $f^{*}:\mathcal{X}\to\Delta^{d}$ of $P$ is defined as
@@ 3.2 Loss Function
The function $\psi\circ\rho_{f}$ is a binary classifier, since $\psi$ controls the range of $\rho_{f}$. In fact, the classifier returns the prediction equal to $+1$ if the covariates $x$ and $x^{\prime}$ coincide. In addition, by Theorem 3 in (Alexander, , 1977), the prediction equals $-1$ if and only if these points are embedded in different vertices on $\Delta^{d}$.
Note that a similar classifier defined with the Mahalanobis distance is proposed in (Jin et al., , 2009). On the other hand, in Definition 7, we employ the Euclidean distance and additionally embed the variables $x$ and $x^{\prime}$ in the regular simplex. Furthermore, Jin et al., (2009) define the binary variable by the information showing whether the supervised labels of two covariates coincide or not. Meanwhile, we use the statistical dependence to define the distribution of $(X,X^{\prime},Y)$, as in (Tsai et al., , 2020).
Hinge loss is often studied in the literature on statistical learning (e.g., (Boser et al., , 1992; Lin, , 2002; Zhang, , 2004; Bartlett et al., , 2006; Lecué, , 2007; Kim et al., , 2021)). In particular, in a setting of conventional binary classification, Kim et al., (2021, Theorem 3.2) study nonparametric estimation of smooth boundaries using empirical risk minimization with hinge loss. Note that it is common in the literature (Kim et al., , 2021; Meyer, , 2023; Imaizumi and Fukumizu, , 2019, 2022) that nonparametric estimation of smooth boundaries is studied with some specific loss functions, such as 0-1 loss in (Meyer, , 2023) and squared loss in (Imaizumi and Fukumizu, , 2019, 2022). Furthermore, hinge loss and its variants are also studied in the context of self-supervised learning (Arora et al., , 2019; Li et al., , 2021; Shah et al., , 2022; Waida et al., , 2023; Ji et al., , 2023) and similarity learning (Jin et al., , 2009; Cao et al., , 2016; Zhou et al., , 2024). We use this loss function in a contrastive learning algorithm, as it belongs to a general class of loss functions of contrastive learning proposed in (Chen et al., 2021b, ).
@@
  • Note that the relationship between regular simplices and empirical risk minimizers in representation learning is shown in (Liu et al., , 2021; Lee et al., , 2024; Koromilas et al., , 2024). On the other hand, in the proposition above we are mainly interested in the population risk minimizers.
  • Zhou et al., (2024, Theorem 1) show that the minimizer of the population risk with respect to a hinge loss defined in a similarity learning setting embeds covariates in the probability simplex. Since Proposition 8 deals with a different problem setting, the assumptions used in the present work are distinct from the ones considered in (Zhou et al., , 2024).
@@ 3.3 Classes of Vector-Valued Functions
Let $\beta>0$, $\beta_{0}\geq 0$, and $P\in\mathcal{P}$. Then, the $(\beta,\beta_{0},P)$-localized subclass of a set $\mathcal{F}\subset\mathcal{F}_{0}$, which is denoted by $\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$, is defined as
@@
Note that some related concepts are introduced in (Schiebinger et al., , 2015; Trillos et al., , 2021) in the context of clustering theory. Indeed, if $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$ for some $\mathcal{F}\subseteq\mathcal{F}_{0}$, $\beta>0$, $\beta_{0}\geq 0$, and $P\in\mathcal{P}$, then it holds that
@@ 4.2.2 A General Estimation Bound
Let $P\in\mathcal{P}$, and let $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ be any sequence in $\mathscr{S}_{P}$. Denote by $f^{*}$, the contrastive function of $P$. Let $\beta\in(0,D_{{\textup{proj}}})$, $\beta_{0}\geq 0$, and $\mathcal{F}\subset\mathcal{F}_{0}$. For any distinct $i,j\in\{1,\cdots,d_{1}\}$ and every $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$, we have
@@
Let $\beta\in(0,D_{{\textup{proj}}})$, $\beta_{0}\geq 0$, $n\in\mathbb{N}\setminus\{1\}$, $\mathcal{F}\subset\mathcal{F}_{0}$, and $P\in\mathcal{P}$. Let $f^{*}$ denote the contrastive function of $P$, and let $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ be any sequence in $\mathscr{S}_{P}$. Then, there exists a constant $C_{2}>0$ such that for any $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$ we have
@@
Let $\beta\in(0,D_{{\textup{proj}}})$, $\beta_{0}\geq 0$, $\mathcal{F}\subset\mathcal{F}_{0}$, and $P\in\mathcal{P}$. Let $f^{*}$ be the contrastive function of $P$. Let $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ be any sequence in $\mathscr{S}_{P}$. For every $i,j\in\{1,\cdots,d_{1}\}$ such that $i\neq j$, any $r\in(0,2^{-1}\beta]$, and any $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$, there exists a constant $C_{i,j}>0$ independent of $r$ such that
@@ 4.3.2 Related Work on Pairwise Binary Classification
In the proof of Theorem 12, we show an inequality between two quantities, namely $\|f-f^{*}\|_{\mathcal{X},P_{X},2}$ and $\mathbb{E}_{P_{X,X^{\prime}}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]$. A similar inequality for a contrastive learning problem has recently been shown in (Ge et al., , 2024). Specifically, Ge et al., (2024) consider the setting where the paired covariates $(X,X^{\prime})$ are independent of each other to show an inequality in (Ge et al., , 2024, Lemma D.2). On the other hand, we need a largely different approach to show our inequality since we consider the problem setting where the covariates $X$ and $X^{\prime}$ are not necessarily independent, following (Tsai et al., , 2020). Since the independence assumption is used in (Ge et al., , 2024, Lemma D.2), our inequality is not shown by their approach. Furthermore, in the field of contrastive learning, it is common to assume that $X$ and $X^{\prime}$ can be dependent, both in practice and in theory (see, e.g, (Arora et al., , 2019; Chen et al., , 2020; HaoChen et al., , 2021)). Therefore, Theorem 12 also gives some insights of how we can address the technical difficulty, which is described with localization of function classes as shown in Section 4.2.2.
@@ 5 Application to Multiclass Classification
    There is some $P\in\mathcal{P}_{\alpha}$ such that the marginal distributions of $P$ and $P_{X,Z_{d_{1}}}$ with respect to $\mathcal{X}$ coincide with each other, and for any partition $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ in $\mathscr{S}_{P}$, any Bayes classifier $z^{*}$ of $P_{X,Z_{d_{1}}}$ satisfies $D_{P_{X}}((z^{*})^{-1}(i),\mathcal{K}_{i})=0$.

### v2

@@ 2.2 Assumptions
We next recall the classes of smooth boundaries introduced by Imaizumi and Fukumizu, (2022). For any $\alpha>0$, $R\geq 0$, and $d\in\mathbb{N}$, denote by $\mathcal{C}^{\alpha,d}$, the ball of the $\alpha$-Hölder space on $[0,1]^{d}$ with radius $R$, namely
$\mathcal{C}^{\alpha,d}=\left\{g:[0,1]^{d}\to\mathbb{R}\;|\;g\textup{ is }\alpha_{*}\textup{-times differentiable, }\|g\|_{\mathcal{C}^{\alpha,d}}\leq R\right\},$
@@
    Given $\mathscr{S}=\{\mathcal{J}_{i}\}_{i=1}^{d_{1}}$, there are disjoint partition $\{\mathcal{I}_{1},\cdots,\mathcal{I}_{d_{1}}\}$ of $\{1,-1\}^{d_{*}}$, functions $h_{1},\cdots,h_{d_{*}}\in\mathcal{C}^{\alpha,d_{0}-1}$, and indices $j_{1},\cdots,j_{d_{*}}\in\{1,\cdots,d_{0}\}$, such that for every $i=1,\cdots,d_{1}$,
@@
    Given $\mathscr{S}^{\prime}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$, there are disjoint partition $\{\mathcal{I}_{1},\cdots,\mathcal{I}_{d_{1}}\}$ of $\{1,-1\}^{d_{*}}$, functions $h_{1},\cdots,h_{d_{*}}\in\mathcal{C}^{\alpha,d_{0}-1}$, and indices $j_{1},\cdots,j_{d_{*}}\in\{1,\cdots,d_{0}\}$ such that for every $i=1,\cdots,d_{1}$,
@@
Definition of the map $\mathscr{S}_{P}$.
Since any partition $\mathscr{S}$ in $\mathscr{P}_{\alpha}$ is expressed as a sequence, we need to introduce an equivalence relation that identifies the sequences containing the same subsets: namely, given two sequences $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ and $\{\mathcal{K}_{i}^{\prime}\}_{i=1}^{d_{1}}$ of subsets in $\mathcal{X}$, we define the equivalence relation $\sim$ as $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\sim\{\mathcal{K}_{i}^{\prime}\}_{i=1}^{d_{1}}$ if and only if there is a permutation $\pi$ of the set $\{1,\cdots,d_{1}\}$ such that $\mathcal{K}_{\pi(i)}=\mathcal{K}_{i}^{\prime}$ for any $i\in\{1,\cdots,d_{1}\}$.
@@
Given $\alpha>0$, we define the map $P\mapsto\mathscr{S}_{P}$ on $\mathcal{P}_{\alpha}$ so that the following condition is satisfied for every $P\in\mathcal{P}_{\alpha}$: for any partition $\mathscr{S}\in\mathscr{P}_{\alpha}$ satisfying the condition (A4) for $P$, we have $\mathscr{S}_{P}\sim\mathscr{S}$.
Note that by the condition (4), the map $P\mapsto\mathscr{S}_{P}$ on $\mathcal{P}_{\alpha}$ is well-defined. Also in the above definition, without loss of generality we can fix an arbitrary partition $\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}\sim\mathscr{S}$ for each $P\in\mathcal{P}_{\alpha}$, by the definition of the equivalence relation.
@@ 2.3 Formal Statement of the Problem
Here, $\|\cdot\|_{1}$ denotes the 1-norm in the Euclidean space. Given $\alpha>0$, let $P\in\mathcal{P}_{\alpha}$. Then, we aim at estimating the sets in the sequence $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$. Similarly to Imaizumi and Fukumizu, (2019, 2022), we analyze the convergence rates of estimators in terms of the $L^{2}$-risk. Formally, the estimation error of an estimator $\widehat{g}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$ is measured by the sum of $L^{2}$-risks
@@
Given $P\in\mathcal{P}_{\alpha}$, is there an estimator $\widehat{g}_{n}:(\mathcal{X}^{2}\times\mathcal{Y})^{n}\to\mathcal{G}_{0}$ such that the $L^{2}$-risk (5) defined with the sequence $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$ tends to zero as $n\to\infty$? In particular:
@@
    Is it reasonable to use the pairwise dependence to estimate smooth boundaries?
@@ 2.4 Validity of the Estimation Problem
For every $\mathscr{S}\in\mathscr{P}_{\alpha,+}$, there is a Borel probability measure $P\in\mathcal{P}_{\alpha}$ such that we have $\mathscr{S}\sim\mathscr{S}_{P}$.
@@ 3.1 Preliminary
Given $P\in\mathcal{P}$, let $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$. The contrastive function $f^{*}:\mathcal{X}\to\Delta^{d}$ of $P$ is defined as
@@ 3.2 Loss Function
The function $\psi\circ\rho_{f}$ is a binary classifier, since $\psi$ controls the range of $\rho_{f}$. In fact, the classifier returns the prediction equal to $+1$ if the vectors $f(x)$ and $f(x^{\prime})$ coincide. In addition, by Theorem 3 in (Alexander, , 1977), the prediction equals $-1$ if and only if these points are embedded in different vertices on $\Delta^{d}$.
We give several comments on Definition 7.
  • A similar classifier defined with the Mahalanobis distance is proposed in (Jin et al., , 2009). On the other hand, in Definition 7, we employ the Euclidean distance and additionally embed the variables $x$ and $x^{\prime}$ in the regular simplex. Furthermore, Jin et al., (2009) define the binary variable by the information showing whether the supervised labels of two covariates coincide or not. Meanwhile, we use the statistical dependence to define the distribution of $(X,X^{\prime},Y)$, as in (Tsai et al., , 2020).
  • Hinge loss is often studied in the literature on statistical learning (e.g., (Boser et al., , 1992; Lin, , 2002; Zhang, , 2004; Bartlett et al., , 2006; Lecué, , 2007; Steinwart and Christmann, , 2008; Kim et al., , 2021)). In particular, in a setting of conventional binary classification, Kim et al., (2021, Theorem 3.2) study nonparametric estimation of smooth boundaries using empirical risk minimization with hinge loss. Note that it is common in the literature (Kim et al., , 2021; Meyer, , 2023; Imaizumi and Fukumizu, , 2019, 2022) that nonparametric estimation of smooth boundaries is studied with some specific loss functions, such as 0-1 loss in (Meyer, , 2023) and squared loss in (Imaizumi and Fukumizu, , 2019, 2022). Furthermore, hinge loss and its variants are also studied in the context of self-supervised learning (Arora et al., , 2019; Li et al., , 2021; Shah et al., , 2022; Waida et al., , 2023; Ji et al., , 2023) and similarity learning (Jin et al., , 2009; Cao et al., , 2016; Zhou et al., , 2024). We use this loss function in a contrastive learning algorithm, as it belongs to a general class of loss functions of contrastive learning proposed in (Chen et al., 2021b, ).
@@
  • The relationship between regular simplices and empirical risk minimizers in representation learning is shown in (Liu et al., , 2021; Lee et al., , 2024; Koromilas et al., , 2024). Meanwhile, in the proposition above we are interested in the population risk minimizers.
  • Zhou et al., (2024, Theorem 1) show that the vector-valued function minimizing the population risk with respect to a hinge loss defined in a similarity learning setting embeds covariates in the probability simplex. Since Proposition 8 deals with a different problem setting, the assumptions used in the present work are distinct from the ones considered in (Zhou et al., , 2024).
@@ 3.3 Classes of Vector-Valued Functions
Let $\beta>0$, $\beta_{0}\geq 0$, and $P\in\mathcal{P}$. Denote by $f^{*}$, the contrastive function of $P$. Then, the $(\beta,\beta_{0},P)$-localized subclass of a set $\mathcal{F}\subset\mathcal{F}_{0}$, which is denoted by $\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$, is defined as
@@
In several different contexts, some similar conditions are studied:
  • In the theory of clustering, some related concepts are introduced in (Schiebinger et al., , 2015; Trillos et al., , 2021). Indeed, if $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$ for some $\mathcal{F}\subseteq\mathcal{F}_{0}$, $\beta>0$, $\beta_{0}\geq 0$, and $P\in\mathcal{P}$, then it holds that
@@ 4.2.2 A General Estimation Bound
Let $P\in\mathcal{P}$, and let $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$. Denote by $f^{*}$, the contrastive function of $P$. Let $\beta\in(0,D_{{\textup{proj}}})$, $\beta_{0}\geq 0$, and $\mathcal{F}\subset\mathcal{F}_{0}$. For any distinct $i,j\in\{1,\cdots,d_{1}\}$ and every $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$, we have
@@
Let $\beta\in(0,D_{{\textup{proj}}})$, $\beta_{0}\geq 0$, $n\in\mathbb{N}\setminus\{1\}$, $\mathcal{F}\subset\mathcal{F}_{0}$, and $P\in\mathcal{P}$. Let $f^{*}$ denote the contrastive function of $P$, and let $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$. Then, there exists a constant $C_{2}>0$ such that for any $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$ we have
@@
Let $\beta\in(0,D_{{\textup{proj}}})$, $\beta_{0}\geq 0$, $\mathcal{F}\subset\mathcal{F}_{0}$, and $P\in\mathcal{P}$. Let $f^{*}$ be the contrastive function of $P$. Let $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$. For every $i,j\in\{1,\cdots,d_{1}\}$ such that $i\neq j$, any $r\in(0,2^{-1}\beta]$, and any $f\in\mathscr{F}_{\beta,\beta_{0},P}(\mathcal{F})$, there exists a constant $C_{i,j}>0$ independent of $r$ such that
@@ 4.3.2 Related Work on Pairwise Binary Classification
In the proof of Theorem 12, we show an inequality between $\|f-f^{*}\|_{\mathcal{X},P_{X},2}^{2}$ and $\mathbb{E}_{P_{X,X^{\prime}}}[|\psi\circ\rho_{f}-\psi\circ\rho_{f^{*}}|]$. A similar inequality for a contrastive learning problem has recently been shown in (Ge et al., , 2024). Specifically, in Lemma D.2 of (Ge et al., , 2024), they prove an upper bound of $\|f-f^{\prime}\|_{\mathcal{X},P_{X},2}^{2}$ based on the assumption that $\mathbb{E}_{P_{X}}[f(x)f^{\prime}(x)^{\top}]$ is a symmetric matrix for some given vector-valued functions $f,f^{\prime}$. However, this is a strong condition in the setting where only pairwise binary classification data is observed. Also, Ge et al., (2024) consider the setting where the paired covariates $X,X^{\prime}$ are independent. On the other hand, in our setting, the covariates $X$ and $X^{\prime}$ are not necessarily independent, following (Tsai et al., , 2020). Note that, in the field of contrastive learning, it is common to assume that $X$ and $X^{\prime}$ can be dependent, both in practice and in theory (see, e.g, (Arora et al., , 2019; Chen et al., , 2020; HaoChen et al., , 2021)). Therefore, Theorem 12 also gives some insights of how we can address the technical difficulties, which is described with localization of function classes as shown in Section 4.2.2.
@@ 5 Application to Multiclass Classification
    There is some $P\in\mathcal{P}_{\alpha}$ such that the marginal distributions of $P$ and $P_{X,Z_{d_{1}}}$ with respect to $\mathcal{X}$ coincide with each other, and for the partition $\mathscr{S}_{P}=\{\mathcal{K}_{i}\}_{i=1}^{d_{1}}$, any Bayes classifier $z^{*}$ of $P_{X,Z_{d_{1}}}$ satisfies $D_{P_{X}}((z^{*})^{-1}(i),\mathcal{K}_{i})=0$.